OpenStudy (micheal1234):

SOLVING EQUATIONS SAMPLE WORK Lesson 1: Introduction to Equations Lesson Vocabulary Equation Open Sentence Solution to an Equation 1. Is 2 a solution of the following equation? 2 = 10 – 4y 2. Use a table to find the solution to 5x + 3 = 23 3. Which equation has a solution of ½? A. 13x – 12 = 14 B. 9x + 15 = 20 C. -6x – 18 = -21 D. -11x = 12x + 12

OpenStudy (misty1212):

HI!!

OpenStudy (micheal1234):

hi

OpenStudy (misty1212):

Is 2 a solution of the following equation? 2 = 10 – 4y ? put a 2 where you see yq

OpenStudy (misty1212):

is \[2-10-4\times 2\]?

OpenStudy (micheal1234):

??

OpenStudy (darrien117):

????

OpenStudy (micheal1234):

i say move the 4y to the other side

OpenStudy (misty1212):

oops

OpenStudy (misty1212):

is \[2=10-4\times 2\]

OpenStudy (micheal1234):

4y+2=10

OpenStudy (misty1212):

there is nothing to "move" it is only asking you to check

OpenStudy (micheal1234):

ok

OpenStudy (misty1212):

true or false?\[2=10-4\times 2\]

OpenStudy (texaschic101):

your just subbing in 2 for y and then seeing if the equation is true

OpenStudy (micheal1234):

so do we 4 ×2 first

OpenStudy (texaschic101):

yes

OpenStudy (micheal1234):

2=10-8=2

OpenStudy (micheal1234):

thank for helping

OpenStudy (micheal1234):

i ready got 2,3